# EMAGE: Towards Unified Holistic Co-Speech Gesture Generation via Expressive Masked Audio Gesture Modeling

### v4

@@ Abstract
We propose EMAGE, a framework to generate full-body human gestures from audio and masked gestures, encompassing facial, local body, hands, and global movements. To achieve this, we first introduce BEAT2 (BEAT-SMPLX-FLAME), a new mesh-level holistic co-speech dataset. BEAT2 combines MoShed SMPLX body with FLAME head parameters and further refines the modeling of head, neck, and finger movements, offering a community-standardized, high-quality 3D motion captured dataset. EMAGE leverages masked body gesture priors during training to boost inference performance. It involves a Masked Audio Gesture Transformer, facilitating joint training on audio-to-gesture generation and masked gesture reconstruction to effectively encode audio and body gesture hints. Encoded body hints from masked gestures are then separately employed to generate facial and body movements. Moreover, EMAGE adaptively merges speech features from the audio’s rhythm and content and utilizes four compositional VQ-VAEs to enhance the results’ fidelity and diversity. Experiments demonstrate that EMAGE generates holistic gestures with state-of-the-art performance and is flexible in accepting predefined spatial-temporal gesture inputs, generating complete, audio-synchronized results. Our code and dataset are available111*Equal Contribution. [^2] https://pantomatrix.github.io/EMAGE/
@@ 1 Introduction
The final stretch in Full-body Co-Speech Gesture Generation is brief yet riddled with challenges. Despite the development of various baselines, e.g., audio conditioned models for face [19, 60, 15], body [65, 25, 13], and hand movements [40, 39, 33], along with a few attempts at merged models [28, 64], the limited availability of comprehensive data and models poses an ongoing obstacle. Instead of creating results that can directly substitute for the delicate designs of animators, in this paper, we consider a framework that can incorporate partial spatial-temporal predefined gestures and autonomously complete the remaining frames in synchronization with the audio, offering advantages in applications for creating photo-realistic digital humans[63, 56, 62, 70].
To achieve this, we first require a comprehensive gesture dataset. While several datasets are available for audio-to-body [65], audio-to-body&face [28], body-to-hands [40, 39], and common actions [45, 16], integrating them poses challenges due to varying data formats and evaluation metrics across different sub-areas. This motivates us to establish a unified, community-standard benchmark for training and evaluating holistic gestures. Most close to this end, although BEAT [39] builds the most comprehensive mocap co-speech gesture dataset with diverse modalities. Despite its size and richness, the data in BEAT is not presented in a standardized format. The dataset comprises iPhone ARKit blendshape weights and Vicon skeletons, posing challenges in training to achieve a satisfied performance due to the absence of a unified representation, e.g., vertices for vertex loss [51, 20]. Additionally, conducting direct human perceptual evaluation on these diverse modalities is inherently complex. On the other hand, SMPLX [49] and FLAME [35] are common mesh standards for the academic community for other human dynamics-related tasks, e.g., action [51] and talking head generation [14]. Driven by the motivation to facilitate the sharing of knowledge across different tasks, we present BEAT-SMPLX-FLAME (BEAT2), which consists of two main components: i) Refined SMPLX body shape and pose parameters from Mosh++ with hard-coded physical priors. ii) Optimized high-quality FLAME head parameters. By integrating SMPLX’s body and FLAME’s head, BEAT2 facilitates a comprehensive training and evaluation benchmark across multiple sub-areas in the context of co-speech human animation generation.
@@
Figure 2: Comparison of Data between BEAT2 and Others. BEAT-SMPLX-FLAME presents a new mesh-level, motion-captured, holistic co-speech gesture dataset with 60h of data. Left: We compare our refined SMPLX body parameters results (denoted as Refined Mosh) with Original BEAT’s Skeleton [39], Retargeted from AutoRegPro, and initial results of Mosh++ [49]. The refined results show correct neck flexion, appropriate head and neck shape ratios, and detailed finger representations. Right: Visualization of blendshape weights from Original BEAT [39] with ARKit’s template, Wrapped based, and handcrafted optimization. Our final handcrafted FLAME blendshape-based optimization demonstrates both accurate lip movement details and natural mouth shapes.
With the holistic data from BEAT2, we aim to enhance the coherence of individual body parts while ensuring accurate cross-modal alignment between audio and motion, c.f. Figure 1. This leads to the design of Expressive Masked Audio-Conditioned Gesture Modeling, denoted as EMAGE, a spatial and temporal transformer-based framework. EMAGE first aggregates spatial features from masked body joints. Then, it adeptly reconstructs the latent space of pretrained gestures with a switchable gesture temporal self-attention and audio-gesture cross-attention. The selection of different forward paths enables modeling the gesture-to-gesture and audio-to-gesture prior separately and effectively. Once the reconstructed latent features are obtained, EMAGE decodes local face and body gestures from four compositional pretrained Vector Quantized-Variational AutoEncoders (VQ-VAEs) and decodes global translations from a pretrained Global Motion Predictor. The compositional VQ-VAEs are key to preserving audio-related motions during the reconstruction. Through these designs, EMAGE achieves state-of-the-art performance in generating body and face gestures. It is flexible to accept audio and partially initialized gestures, recovering audio-synchronized, coherent gestures. More excitingly, EMAGE showcases its further applications in joint training multiple, non-holistic gesture datasets.
Overall, our contributions are as follows:
  • We release BEAT2, a representation-unified, mesh-level dataset that leverages Moshed SMPLX-body and FLAME head parameters.
  • We propose EMAGE, a simple yet effective holistic gesture generation framework that generates coherent gestures with partial gesture and audio priors.
  • As same as previous methods, EMAGE achieves state-of-the-art performance in generating both body and face gestures with only four frames of seed gestures.
  • EMAGE showcases the use of additional non-holistic gesture datasets for training, e.g., Trinity and AMASS, to further improve the fidelity and diversity of the results, effectively leveraging data from different datasets.
@@ 2 Related Work
Co-Speech Animation Datasets are categorized into two types, c.f. Table 1: pseudo-labeled (PGT) and motion-captured (mocap). For PGT, the Speech2Gesture Dataset [25] utilizes OpenPose [11] to extract 2D poses from News and Teaching videos. Subsequent works extend the dataset to 3D pose [28] and SMPLX [64]. The TED Dataset [66], which extracts 2D pose from TED videos, has been extended to include 3D pose [65], fingers [41], and SMPLX [44] in subsequent years. Similarly, 2D, 3D landmarks, and mesh are estimated from multi-view recorded videos [59] as PGT for talking head generation. Although pseudo-labeled approaches could theoretically extract infinite data, their accuracy is limited. For instance, the average error (in mm) for the SOTA monocular 3D pose estimation algorithm [23] in the Human 3.6M dataset [30] is 33.4, while for Vicon mocap it is 0.142 [46]. For mocap datasets, Trinity [3] features one male actor with 4h of data, and TalkingWithHands [32] collects conversation scenarios for two speakers. ZEEG [24] considers one speaker with 12 styles. For the face, 3D scanning is accurate but limited in quantity due to cost, e.g., less than 3h for BIWI [21], VOCA [14], and MeshTalk [55]. This leads to a balance between performance and quantity in the ARKit dataset [50]. The aforementioned datasets address the body and face separately. On the other hand, BEAT [39] includes both 3D pose body and ARKit [8] facial blendshape weights for the first time. However, BEAT lacks mesh data.
@@
Co-Speech Animation Models are categorized based on their outputs, c.f. Table 2. In addition to the baselines and datasets [25, 65, 14, 55, 32, 47], several frameworks for body gestures [52, 53, 2, 33] have improved the performance of the baselines. They are trained and evaluated by selecting specific upper body joints [36, 39, 4, 5, 41, 69, 48], or all body joints [71, 61]. Recent improvements in facial gestures drive the vertex by transformer and discrete face priors [20, 60]. However, these methods only address either the face or body. Applying these methods to the full body will yield sub-optimal results as audio has different correlations with face and body dynamics. Most similar to our work, Habibie et al. [28] employs a single audio encoder and multiple decoders for generating face and body gestures. Talkshow [64] demonstrates the advantage of separating the audio encoders and decoding the body and hands using quantized codes in an auto-regressive manner. However, it lacks lower body and global motion, and there is no shared information for body and facial gesture generation. Moreover, it cannot accept partially masked body hints due to the design of a fully auto-regressive model.
@@
Masked Representation Learning was first demonstrated to be effective in Natural Language Processing, with Bert-based models [17, 31, 42] boosting the performance of learned word embeddings in downstream tasks through a combination of masked language modeling and transformer architecture. Subsequently, Masked AutoEncoders [29] expanded masked image modeling to computer vision by removing inputted image patches. This concept of masked representation learning has since been employed in other modalities, e.g., video [18, 12, 43], audio [26, 37, 38], and point cloud [68, 67, 54]. Most related to our work, MotionBERT [72] proposes a spatial-temporal transformer to learn robust motion representation for classification-based tasks by masking 2D pose. Differing from their method, we target robust motion features for conditional motion generation tasks, requiring a balance in training between multiple modalities, e.g., audio and gesture.
@@ 3 BEAT2
We introduce how to obtain unified mesh-level data, i.e., SMPLX [49] and FLAME [35] parameters, from the original BEAT [39] in this section. BEAT utilizes a Vicon motion capture system [46] and releases 78-joint skeleton-level Biovision Hierarchy (BVH) [10] motion files. Their facial capture system uses ARKit [8] with a depth camera on the iPhone 12 Pro, extracting $\mathbb{R}^{51}$ blendshape weights. These blendshapes are designed based on the Facial Action Coding System (FACS) and are widely adopted by industry novices. However, both the motion and facial data lack mesh-level details (c.f. Figure 2), e.g., shapes and vertices.
@@ 3.1 Body Parameters Initialization via Mosh++
We initialize the SMPLX body shape and pose parameters from BEAT captured markers data based on Mosh++ [49]. Given the captured markers positions $\mathbf{m}\in\mathbb{R}^{T\times K\times 3}$, predefined markers positions offset $\mathbf{d}\in\mathbb{R}^{K\times 3}$, and user-defined vertices to marker function $\mathcal{H}$, we aim to obtain body shape $\beta\in\mathbb{R}^{300}$, poses $\theta\in\mathbb{R}^{T\times 55\times 3}$, and translation parameters $\gamma\in\mathbb{R}^{T\times 3}$. The optimization uses differentiable surface vertices mapping function $\mathcal{S}(\beta,\theta,\gamma)$ and vertex normal function $\mathcal{N}(\beta,\theta_{0},\gamma_{0})$. For each frame, the latent marker $\tilde{\mathbf{m}}\in\mathbb{R}^{T\times K\times 3}$ is calculated as
@@
We first select 12 frames to optimize and fix $\mathbf{d}$ and $\beta$, then optimize $\theta_{i}$, $\gamma_{i}$ for $i\in(1:T)$ by minimizing the loss terms based on $\|\tilde{\mathbf{m}}_{i}-\mathbf{m}_{i}\|^{2}$, including Data Term, Surface Distance Energy, Marker Initialization Regularization, Pose and Shape Priors, Velocity Constancy, and Soft-Tissue Term (cf. supplementary materials). The overall objective function is the weighted sum of these terms to balance accuracy and plausibility.
@@ 3.2 Body Parameters Refinement
We refine body shape and pose parameters with three simple yet effective hard-coded physical rules. i) The neck and head length should approximate $1/7$ of the total length of the body. ii) Fingers, except for the thumb, should not bend backward. iii) We employ the Kolmogorov-Smirnov test, which reveals the data is similar to a normal distribution. Subsequently, we apply a Gaussian truncation approach, where all data points falling outside the $3\sigma$ range are adjusted to align with the $3\sigma$ threshold and blended with adjacent 10 frames. We compare our refined body parameters (denoted as Mosh Refined) with Original BEAT, Retargeted SMPLX, and Mosh SMPLX in Figure 2.
@@ 4 EMAGE
We introduce the details of EMAGE, Expressive Masked Audio-Conditioned Gesture Modeling (c.f. Figure 3). Given gestures $\mathbf{g}\in\mathbb{R}^{T\times(55\times 6+100+4+3)}$, representing 55 joints in Rot6D, $\mathbb{R}^{100}$ FLAME parameters, $\mathbb{R}^{4}$ foot contact labels, $\mathbb{R}^{3}$ global translations, and speech audio $\mathbf{a}\in\mathbb{R}^{T\times sk}$, where $sk=sr_{\text{audio}}/fps_{\text{gestures}}$, EMAGE jointly optimizes masked gesture reconstruction and audio-conditioned gesture generation. This optimization enhances performance during inference and enables the use of partially masked gestures to complete holistic gestures. To achieve this, we first model the quantized latent space in Sec. 4.1, following [64, 60]. We then design a separated speech audio encoder using Content Rhythm Self-Attention (CRA) as detailed in Sec. 4.2. Subsequently, we learn the body hints from masked gestures via the Masked Audio Gesture Transformer, explained in Sec. 4.3. Finally, the decoding strategy for different body segments is discussed in Sec. 4.4.

### v5

@@ Abstract
We propose EMAGE, a framework to generate full-body human gestures from audio and masked gestures, encompassing facial, local body, hands, and global movements. To achieve this, we first introduce BEAT2 (BEAT-SMPLX-FLAME), a new mesh-level holistic co-speech dataset. BEAT2 combines a MoShed SMPL-X body with FLAME head parameters and further refines the modeling of head, neck, and finger movements, offering a community-standardized, high-quality 3D motion captured dataset. EMAGE leverages masked body gesture priors during training to boost inference performance. It involves a Masked Audio Gesture Transformer, facilitating joint training on audio-to-gesture generation and masked gesture reconstruction to effectively encode audio and body gesture hints. Encoded body hints from masked gestures are then separately employed to generate facial and body movements. Moreover, EMAGE adaptively merges speech features from the audio’s rhythm and content and utilizes four compositional VQ-VAEs to enhance the results’ fidelity and diversity. Experiments demonstrate that EMAGE generates holistic gestures with state-of-the-art performance and is flexible in accepting predefined spatial-temporal gesture inputs, generating complete, audio-synchronized results. Our code and dataset are available.*Equal Contribution. [^1] https://pantomatrix.github.io/EMAGE/
@@ 1 Introduction
The path towards full-body co-speech gesture generation presents many remaining challenges. Despite the development of various baselines, e.g., audio conditioned models for face [20, 61, 15], body [67, 25, 13], and hand movements [41, 39, 33], along with a few attempts at merged models [28, 65], the limited availability of comprehensive data and models poses an ongoing obstacle. To make progress in this direction, we present a new framework that can incorporate partial spatial-temporal predefined gestures and autonomously complete the remaining frames in synchronization with the audio. This provides a new tool for creating realistic animations of digital humans [64, 57, 63, 71].
To achieve this, we first require a comprehensive gesture dataset. While several datasets are available for audio-to-body [67], audio-to-body&face [28], body-to-hands [41, 39], and common actions [46, 16], integrating them poses challenges due to varying data formats and evaluation metrics across different sub-areas. This motivates us to establish a unified, community-standard benchmark for training and evaluating holistic gestures. We build upon the BEAT dataset [39], which provides the most comprehensive mocap co-speech gesture dataset with diverse modalities. Despite its size and richness, the data in BEAT is not presented in a standardized format. The dataset comprises iPhone ARKit blendshape weights and Vicon skeletons, posing challenges when using the data for training; e.g., the lack of a mesh representation prevents the use of a vertex loss [52, 19]. Additionally, conducting direct human perceptual evaluations on these diverse modalities is inherently complex. On the other hand, SMPL-X [50] and FLAME [35] are common mesh standards for the academic community for other human dynamics-related tasks, e.g., action [52] and talking head generation [14]. Driven by the motivation to facilitate the sharing of knowledge across different tasks, we present BEAT-SMPLX-FLAME (BEAT2), which consists of two main components: i) Refined SMPL-X body shape and pose parameters from MoSh++ [44, 46] with hard-coded physical priors. ii) Optimized high-quality FLAME head parameters. By integrating SMPL-X’s body and FLAME’s head, BEAT2 facilitates a comprehensive training and evaluation benchmark across multiple sub-areas of co-speech human animation generation.
@@
Figure 2: Comparison of Data between BEAT2 and Others. BEAT-SMPLX-FLAME presents a new mesh-level, motion-captured, holistic co-speech gesture dataset with 60h of data. Left: We compare our refined SMPL-X body parameters (denoted as Refined MoSh) with the original BEAT skeleton [39], Retargeted from AutoRegPro, and initial results of Mosh++ [50]. The refined results show correct neck flexion, appropriate head and neck shape ratios, and detailed finger representations. Right: Visualization of blendshape weights from the original BEAT dataset [39] with ARKit’s template, Wrapped-based, and handcrafted optimization. Our final handcrafted FLAME blendshape-based optimization demonstrates both accurate lip movement details and natural mouth shapes.
With the holistic data from BEAT2, we aim to enhance the coherence of individual body parts while ensuring accurate cross-modal alignment between audio and motion; see Figure 1. This leads to the design of Expressive Masked Audio-conditioned GEsture modeling, denoted as EMAGE, a spatial and temporal transformer-based framework. EMAGE first aggregates spatial features from masked body joints. Then, it reconstructs the latent space of pretrained gestures with a switchable gesture temporal self-attention and audio-gesture cross-attention. The selection of different forward paths enables modeling the gesture-to-gesture and audio-to-gesture prior separately and effectively. Once the reconstructed latent features are obtained, EMAGE decodes local face and body gestures from four compositional pretrained Vector Quantized-Variational AutoEncoders (VQ-VAEs) and decodes global translations from a pretrained Global Motion Predictor. The compositional VQ-VAEs are key to preserving audio-related motions during the reconstruction. Through these designs, EMAGE achieves state-of-the-art performance in generating body and face gestures. It takes as input audio and partially initialized gestures, and recoveres audio-synchronized, coherent gestures. Additionally, we show how EMAGE can flexibly incorporate additional, non-holistic, datasets to improve results.
Overall, our contributions are as follows: (1) We release BEAT2, a representation-unified, mesh-level dataset that leverages MoShed SMPL-X body and FLAME head parameters. (2) We propose EMAGE, a simple yet effective holistic gesture generation framework that generates coherent gestures with partial gesture and audio priors. (3) EMAGE achieves state-of-the-art performance in generating both body and face gestures with only four frames of seed gestures. (4) EMAGE showcases the use of additional non-holistic gesture datasets for training, e.g., Trinity and AMASS, to further improve the fidelity and diversity of the results, effectively leveraging data from different datasets.
@@ 2 Related Work
Co-Speech Animation Datasets are categorized into two types; see Table 1: pseudo-labeled (PGT) and motion-captured (mocap). For PGT, the Speech2Gesture Dataset [25] utilizes OpenPose [11] to extract 2D poses from News and Teaching videos. Subsequent works extend the dataset to 3D pose [28] and SMPL-X [65]. The TED Dataset [66], which extracts 2D pose from TED videos, has been extended to include 3D pose [67], fingers [40], and SMPL-X [45]. Similarly, 2D, 3D landmarks, and meshes are estimated from multi-view recorded videos [60] as PGT for talking-head generation. Although pseudo-labeled approaches could theoretically extract infinite data, their accuracy is limited. For instance, the average error (in mm) for the SOTA monocular 3D pose estimation algorithm [23] on the Human3.6M dataset [30] is 33.4, while for Vicon mocap it is 0.142 [47]. For mocap datasets, Trinity [3] features one male actor with 4h of data, and TalkingWithHands [32] collects conversation scenarios for two speakers. ZEEG [24] considers one speaker with 12 styles. For the face, 3D scanning is accurate but limited in quantity due to cost, e.g., less than 3h for BIWI [21], VOCA [14], and MeshTalk [56]. This leads to a balance between performance and quantity in the ARKit dataset [51]. The aforementioned datasets address the body and face separately. On the other hand, BEAT [39] includes both 3D pose body and ARKit [8] facial blendshape weights for the first time. However, BEAT lacks mesh data.
@@
Co-Speech Animation Models are categorized based on their outputs; see Table 2. In addition to the baselines and datasets [25, 67, 14, 56, 32, 48], several frameworks for body gestures [53, 54, 2, 33] have improved the performance of the baselines. They are trained and evaluated by selecting specific upper body joints [38, 39, 5, 4, 40, 70, 49], or all body joints [72, 62]. Recent improvements in facial gestures drive the vertices with transformers and discrete face priors [19, 61]. However, these methods only address either the face or body. Applying these methods to the full body will yield sub-optimal results as audio is differently correlated with face and body dynamics. Most similar to our work, Habibie et al. [28] employ a single audio encoder and multiple decoders for generating face and body gestures. TalkSHOW [65] demonstrates the advantage of separating the audio encoders and decoding the body and hands using quantized codes in an auto-regressive manner. However, it lacks lower body and global motion, and there is no shared information for body and facial gesture generation. Moreover, it cannot accept partially masked body hints due to the design of a fully auto-regressive model.
@@
Masked Representation Learning was first demonstrated to be effective in Natural Language Processing, with BERT-based models [17, 31, 42] boosting the performance of learned word embeddings in downstream tasks through a combination of masked language modeling and transformer architecture. Subsequently, Masked AutoEncoders [29] expanded masked image modeling to computer vision by removing and inpainting image patches. This concept of masked representation learning has since been employed in other modalities, e.g., video [18, 12, 43], audio [26, 36, 37], and point clouds [69, 68, 55]. Most related to our work, MotionBERT [73] proposes a spatial-temporal transformer to learn a robust motion representation for classification-based tasks by masking 2D pose. In contrast to their method, we target robust motion features for conditional motion generation tasks, requiring a balance in training between multiple modalities, e.g., audio and gesture.
@@ 3 BEAT2
This section introduces how we obtain unified mesh-level data, i.e., SMPL-X [50] and FLAME [35] parameters, from the original BEAT dataset [39]. BEAT utilizes a Vicon motion capture system [47] and releases 78-joint skeleton-level Biovision Hierarchy (BVH) [10] motion files. Their facial capture system uses ARKit [8] with a depth camera on the iPhone 12 Pro, extracting $\mathbb{R}^{51}$ blendshape weights. These blendshapes are designed based on the Facial Action Coding System (FACS), which is widely adopted. However, both the motion and facial data lack mesh-level details (see Figure 2), e.g., shapes and vertices.
@@ 3.1 Body Parameters Initialization via MoSh++
We initialize the SMPL-X body shape and pose parameters from BEAT mocap marker data using MoSh++ [46]. Given the captured markers positions $\mathbf{m}\in\mathbb{R}^{T\times K\times 3}$, predefined markers position offsets $\mathbf{d}\in\mathbb{R}^{K\times 3}$, and user-defined vertices-to-markers function $\mathcal{H}$, we aim to obtain body shape $\beta\in\mathbb{R}^{300}$, poses $\theta\in\mathbb{R}^{T\times 55\times 3}$, and translation parameters $\gamma\in\mathbb{R}^{T\times 3}$. The optimization uses a differentiable surface vertex mapping function $\mathcal{S}(\beta,\theta,\gamma)$ and vertex normal function $\mathcal{N}(\beta,\theta_{0},\gamma_{0})$. For each frame, the latent marker $\tilde{\mathbf{m}}\in\mathbb{R}^{T\times K\times 3}$ is calculated as
@@
We first select 12 frames to optimize and fix $\mathbf{d}$ and $\beta$, then optimize $\theta_{i}$, $\gamma_{i}$ for $i\in(1:T)$ by minimizing the loss terms based on $\|\tilde{\mathbf{m}}_{i}-\mathbf{m}_{i}\|^{2}$, including Data Term, Surface Distance Energy, Marker Initialization Regularization, Pose and Shape Priors, Velocity Constancy, and Soft-Tissue Term (see supplementary materials). The overall objective function is the weighted sum of these terms to balance accuracy and plausibility.
@@ 3.2 Body Parameters Refinement
MoSh++ produces unnatural head shapes due to the fact that head markers were worn on a helmet and it sometimes produces unnatural finger poses. Consequently, we refine body shape and pose parameters with three simple yet effective hard-coded physical rules. i) The neck and head length should approximate $1/7$ of the total length of the body. ii) Fingers, except for the thumb, should not bend backward. iii) We employ the Kolmogorov-Smirnov test, which reveals the data is similar to a normal distribution. Subsequently, we apply a Gaussian truncation approach, where all data points falling outside the $3\sigma$ range are adjusted to align with the $3\sigma$ threshold and blended with the adjacent 10 frames. We compare our refined body parameters (denoted as MoSh Refined) with Original BEAT, Retargeted SMPL-X, and MoSh SMPL-X in Figure 2.
@@ 4 EMAGE
We introduce the details of EMAGE, Expressive Masked Audio-Conditioned GEsture Modeling (see Figure 3). Given gestures $\mathbf{g}\in\mathbb{R}^{T\times(55\times 6+100+4+3)}$, representing 55 joints in Rot6D, $\mathbb{R}^{100}$ FLAME parameters, $\mathbb{R}^{4}$ foot contact labels, $\mathbb{R}^{3}$ global translations, and speech audio $\mathbf{a}\in\mathbb{R}^{T\times sk}$, where $sk=sr_{\text{audio}}/fps_{\text{gestures}}$, EMAGE jointly optimizes masked gesture reconstruction and audio-conditioned gesture generation. This optimization enhances performance during inference and enables the use of partially masked gestures to complete holistic gestures. To achieve this, we first model the quantized latent space in Sec. 4.1, following [65, 61]. We then design a separated speech audio encoder using Content Rhythm Self-Attention (CRA) as detailed in Sec. 4.2. Subsequently, we learn the body hints from masked gestures via the Masked Audio Gesture Transformer, explained in Sec. 4.3. Finally, the decoding strategy for different body segments is discussed in Sec. 4.4.